# Efficient Neuro-Symbolic Learning of Constraints & Objective

### v1

@@ 5.4.2. Training and evaluation
To train the neural network, we use the data set of (Ingraham et al., 2019), already split into train/validation/test sets of respectively $17,000$, $600$, and $1,200$ proteins, in such a way that proteins with similar structures or sequences are in the same set. Similarly to Sudoku, a protein is described by features $\omega$ computed on each pair of positions $i,j$ in the protein. They include inter-atomic distance features encoded with Gaussian radial basis function (Dauparas et al., 2022), and a positional encoding of the sequence distance (|ij||i-j|(Vaswani et al., 2017). Each backbone geometry $\omega$ is associated with $\mathbf{y}$, the sequence of the corresponding known protein. All pair features in $\omega$ are processed by a neural network composed of a recurrent gated MLP (Liu et al., 2021) that learns an environment embedding from a central amino acid and its neighbours within 1010Å, fed to a ResMLP (as for Sudoku) that takes pairs of features and environment embeddings to predict $20\times 20$ cost matrices. Training and architecture details are in Annex C.3.
We train the same model with the same initialization using either the NPLL or the E-PLL loss. To adapt to variable protein sizes, the E-PLL eliminates kk% of incoming neighbor messages. With up to $500$ amino acids, the optimization task is challenging at inference and we used a recent GM convex solver (Durante et al., 2022). One usual metric for protein design is the Hamming distance between the predicted and observed (native) sequences, called the Native Sequence Recovery rate (NSR). Protein design is a multiple-solution problem: above 30% of similarity, two protein sequences are considered as having the same fold (geometry). So, one given structure can adopt many sequences and a 100% NSR cannot be reached.
@@
As shown in Table 4, we observe that the E-PLL not only preserves the good properties of the NPLL but actually improves the NSR. While protein design is often stated as an unconstrained optimization problem, we hypothesize that this improvement results from the existence of infeasibilities: when local environments are very tight, they absolutely forbid large amino acids. Such infeasibilities could not be properly estimated by the NPLL alone.
Our architecture provides a decomposable scoring function, such as those used for protein design in Rosetta (Park et al., 2016). We compared both approaches on the dataset of (Park et al., 2016) in Table 5. Note that small proteins tend to be less constrained than larger ones, hence the decrease in NSR (Ingraham et al., 2019) compared to Table 4. The cost functions learned with our E-PLL outperforms Rosetta’s energy function, leading to designed proteins more similar to natural ones. This is remarkable as Rosetta’s full-atom function considers all atoms of the protein while we just use the backbone geometry and amino acids identities (as in coarse-grained scoring functions (Kmiecik et al., 2016)).
Rosetta [^1] Rosetta’s results are extracted from (Ingraham et al., 2019) E-PLL
@@
Table 5. Comparison with the energy-based design method Rosetta on small single-chain proteins
@@ 6. Conclusion
In this paper, we introduced a neuro-symbolic architecture and a dedicated loss function for learning how to solve discrete reasoning and optimization problems. The architecture is differentiable and as such, allows natural inputs to participate in the definition of discrete reasoning/optimization problems, providing the ability to inject suitable inductive biases that can also enhance data efficiency, as shown in the case of Sudoku. While optimization layers (Pogančić et al., 2020; Sahoo et al., 2023; Wang et al., 2019) can be inserted in an arbitrary position in a neural net, our neuro-symbolic layer with the E-PLL loss must be the final layer. This restriction is of no consequence in many practical settings, where the discrete decision is the final expected output. It instead offers scalable training, avoiding calls to exact solvers that quickly struggle with the noisy instances that are predicted in early training epochs. It is able to benefit from exact or relaxed solvers during inference. Thanks to the E-PLL, it can simultaneously identify an objective to optimize and constraints to satisfy. Finally, its output can be scrutinized to check properties and can be a posteriori completed with side constraints or additional objectives, to inject further instance-dependent information, that may have been independently learned, be available as knowledge, or as user requirements. These facilities are crucially needed in applicative contexts such as protein design (Albanese et al., 2025).
Much remains to be done around this architecture. As it has been done for SAT-Net (Lim et al., 2022), the ultimate $N(\omega)$ GM layer of our architecture could be analyzed during training to identify emerging hypothetical global properties such as symmetries or global decomposable constraints, allowing for more interpretable results.
For memory and computational efficiency reasons, we restricted ourselves to pairwise GMs, limiting the detection of complex many-body interactions. The ability of the architecture to deal with latent variables could be combined with dual/hidden representations of constraints (Stergiou and Walsh, 1999) to enhance its capacity to fully capture such interactions.
The language of NP-hard discrete Graphical Models with latent variables that we use as the output of our architecture is powerful. Yet, like any NP-hard modeling language, it is not necessarily optimal for concisely representing arbitrary decision NP-complete problems. To extend the range of practical problems that can be efficiently and concisely learned by our architecture, the E-PLL loss should be extended with the ability to deal with extended languages, including, e.g., clauses, linear constraints, and other global constraints and cost functions (Bessiere et al., 2023).
@@ Acknowledgements.
We thank S. Buchet and H. Bierlee for proofreading. This work has been financially supported by the French national research project GMLaS under grant agreement ANR-24-CE23-3429 and has been supported by the AI Interdisciplinary Institute ANITI. ANITI is funded by the France 2030 program under the Grant agreement ANR-23-IACL-0002.
@@ B.2. Analysing the PLL as a Fenchel-Young Loss
Fenchel-Young Losses (Blondel et al., 2020) (FYL) are losses that are built based on a regularized prediction function. In the context of SOP, when the structure $\mathcal{Y}$ is captured by a polytope and the objective function is linear, one vertex of the polytope will be an optimal solution. During learning, the learned objective parameters $\mathbf{c}$ evolve, and this leads to sudden shifts of the optimal discrete solution to a new vertex. Regularization in the convex hull $conv(\mathcal{Y})$ of the solution space $\mathcal{Y}$ provides a smoothing component that creates differentiability.
@@
For pairwise MRFs, the sufficient statistics $\phi(\mathbf{y})$ of an assignment $\mathbf{y}$ are defined over pairs of values of pairs of variables $(y_{i},y_{j})$. With the MRF negative log-likelihood (NLL) loss, also known as the Conditional Random Field (CRF) loss, the potential function is the intractable log-partition function (See (Blondel et al., 2020, p. 10)). For tractability reasons, the NPLL is sparse. Its support includes only a neighborhood $N_{1}(\mathbf{y})\subset\mathcal{Y}$ around the true output $\mathbf{y}$, obtained by changing the value of at most one variable:
@@ 0 $100$ - 0%
Table 6. Average performances over $10$ initialization, for various values of parameter $k$. Training is up to $100$ epochs.
@@ C.3. Architecture and training details
For the protein design task, the neural network is composed of an input linear layer of size 128, a block to extract environment information, and a ResMLP to predict cost matrices. The environment network is composed of a gated MLP and a ResNet, repeated $6$ times (with shared weights), the output of the previous iteration being added to the input. The gatedMLP has $3$ layers, a width of $2\times 256$ and it considers the 48 nearest neighbours of each residue. All ResMLP contains $6$ layers, with a recurrent connection every 2 layers, and a width of $256$. The resulting architectures has 3.43.4M parameters.
As in most protein energy functions (Alford et al., 2017), amino acid pairs separated by a large distance are ignored. We use a 15 Å threshold. The neural network is trained using the Adam optimizer, with a weight decay of $10^{-3}$ and an initial learning rate of $5.10^{-4}$, divided by 10 when the validation loss decreases (with patience 0) until it reaches $10^{-*}$. A L1 regularization of $10^{-4}$ is applied to costs. To compare the NPLL loss and the E-PLL loss, we trained the same model with the same hyperparameters and starting from the same weight initialization with each of the loss. Both models run as long as the minimum LR is not reached, and therefore not necessarily for the same number of epochs.

### v3

@@ 5.4.2. Training and Evaluation
To train the neural network, we use the data set of (Ingraham2019, 30), already split into train/validation/test sets of respectively $17,000$, $600$, and $1,200$ proteins, in such a way that proteins with similar structures or sequences are in the same set. Similarly to Sudoku, a protein is described by features $\omega$ computed on each pair of positions $i,j$ in the protein. They include inter-atomic distance features encoded with Gaussian radial basis function (ProteinMPNN, 18), and a positional encoding of the sequence distance (|ij||i-j|(Vaswani2017, 68). Each backbone geometry $\omega$ is associated with $\mathbf{y}$, the sequence of the corresponding known protein. All pair features in $\omega$ are processed by a neural network composed of a recurrent gated MLP (liu2021pay, 41) that learns an environment embedding from a central amino acid and its neighbours within 1010Å, fed to a ResMLP (as for Sudoku) that takes pairs of features and environment embeddings to predict $20\times 20$ cost matrices. Training and architecture details are in Annex C.3.
We train the same model with the same initialization using either the NPLL or the E-PLL loss. To adapt to variable protein sizes, the E-PLL eliminates kk% of incoming neighbor messages. With up to $500$ amino acids, the optimisation task is challenging at inference and we used a recent GM convex solver (durante:hal-03673535, 22). One usual metric for protein design is the Hamming distance between the predicted and observed (native) sequences, called the Native Sequence Recovery rate (NSR). Protein design is a multiple-solution problem: above 30% of similarity, two protein sequences are considered as having the same fold (geometry). So, one given structure can adopt many sequences and a 100% NSR cannot be reached.
@@
As shown in Table 4, we observe that the E-PLL not only preserves the good properties of the NPLL but actually improves the NSR. While protein design is often stated as an unconstrained optimisation problem, we hypothesize that this improvement results from the existence of infeasibilities: when local environments are very tight, they absolutely forbid large amino acids. Such infeasibilities could not be properly estimated by the NPLL alone.
Our architecture provides a decomposable scoring function, such as those used for protein design in Rosetta (RosettaDecoy, 55). We compared both approaches on the dataset of (RosettaDecoy, 55) in Table 5. The dataset is restricted to small proteins because designing with Rosetta on larger proteins is intractable. Note that small proteins tend to be less constrained than larger ones, hence the decrease in NSR (Ingraham2019, 30) compared to Table 4. The cost functions learned with our E-PLL outperforms Rosetta’s energy function, leading to designed proteins more similar to natural ones. This is remarkable as Rosetta’s full-atom function considers all atoms of the protein while we just use the backbone geometry and amino acids identities (as in coarse-grained scoring functions (coarse-grain, 32)).
Table 5. Comparison with the energy-based design method Rosetta on small single-chain proteins.
Rosetta [^1] Rosetta’s results are extracted from (Ingraham2019, 30) E-PLL
@@
6. Discussion
The results we obtained on all variants of the Sudoku-related tasks show impressive data-efficiency compared to pure DL approaches, and it is tempting to try to explain this difference. The first reason that may explain this data efficiency lies in its inductive biases:
  • We focus on learning pairwise interactions and these are exactly those that are present in Sudoku.
  • We take into account the geometry of the grid (cell coordinates being provided as input to the network) and this is crucial for the existence of Sudoku constraints.
However, the first bias is also present in all pure DL-based Transformer-based architectures, as attention matrices are precisely aimed at detecting the same pairwise interactions as in pairwise Graphical Models (bhattacharya2021interpreting, 11). This is also the case for GNN (or RRN) based approaches that focus on (pairwise) graph edges between variables. The input grid-geometry is also provided in Transformer-based approaches, at least indirectly through each variable ROPE positional encoding (wang2025hierarchical, 69).
A more convincing explanation is that learning a compact set of rules should be much simpler than learning a function that maps a partially filled Sudoku grid to a solved grid. Very informally, the space of such functions is huge: considering that an “empty cell” defines an extra state, there are $10^{81}$ possible input grids, and $9^{81}$ possible output grids, thus $90^{81}\approx 10^{158}$ such functions. Instead, the number of pairwise Boolean function networks is $\frac{n(n-1)}{2}\cdot 2^{d^{2}}=81\times 40\times 2^{81}\approx 10^{27}$, which is amazingly smaller. Moreover, contrarily to the function mapping input grids to solutions, the set of rules should be sparse, just because humans don’t like games with complex rules (this also applies to protein design, where a linear number of short-range interactions are known to dominate in physics). In our architecture, L1 regularization directs training towards predicting sparse sets of rules. Eventually, the strategy of learning rules and then exploiting them to find solutions seems therefore far more attractive than directly trying to learn a mapping from input grids to solutions.
Existing pure-DL approaches based on GNNs and Transformers likely follow this path too, but they face a double challenge:
  • They have to learn the correct rules
  • They have to learn how to use them (how to reason on them, using, e.g., Message Passing in GNNs)
We suspect that pure DL architectures struggle on the second challenge. Sudoku is NP-complete and extremely hard grids require more compute than what usual DL models can spend. Iterative denoising diffusion probabilistic models and related hierarchical Transformer-based models (ye2025beyond, 75, 69) can spend more compute at inference and show better results on Sudoku than direct GNN or Transformer-based models (RRN, 54, 73), but they are unable to compete with SOTA symbolic solvers that went through decades of improvements. Although deep DL architectures can emulate dynamic programing (and make it differentiable (mensch2018differentiable, 47)), strongly NP-complete problem solving may either require worst-case exponential space/time dynamic programming (as resolution does in logic (dechter1994directional, 20)) or worst-case exponential time/polytime tree-search (cooper2020graphical, 16). SOTA symbolic solvers exploit a complex combination of both (cooper2020graphical, 16, 46). In contrast, usual DL architectures use fixed space/time or linear (quadratic with Transformers) space/time in denoising/recurrent architectures. Finally, it is not clear how DL architectures can emulate“backtracking”. Even if LLM “reasoning models” could, in principle, do this through their “reasoning trace”, this may require an exponentially long context (something that only very recent Markovian LLM (aghajohari2025markovian, 1) may be able to deal with, considering backtracking requires linear space).
In some sense, the situation here is closely related to symbolic regression (makke2024interpretable, 43) or hybrid genetic programming/MML FunSearch-like approaches (romera2024mathematical, 59): they learn to generate formulae or Python codes that compute a function rather than learning the function directly. In these cases, one can assume the code to be small, which helps learning. Such approaches, however, require assessing the code output during training, which, for NP-complete problem solving, will eventually become excruciating. Our hybrid architecture and its dedicated loss avoid this cost.
@@ 7. Conclusion
In this paper, we introduced a neuro-symbolic architecture and a dedicated loss function for learning how to solve discrete reasoning and optimisation problems. The architecture is differentiable and as such, allows natural inputs to participate in the definition of discrete reasoning/optimisation problems, providing the ability to inject suitable inductive biases that can also enhance data efficiency, as shown in the case of Sudoku. While optimisation layers (blackbox, 57, 60, 70) can be inserted in an arbitrary position in a neural net, our neuro-symbolic layer with the E-PLL loss must be the final layer. This restriction is of no consequence in many practical settings, where the discrete decision is the final expected output. It instead offers scalable training, avoiding calls to exact solvers that quickly struggle with the noisy instances that are predicted in early training epochs. It is able to benefit from exact or relaxed solvers during inference. Thanks to the E-PLL, it can simultaneously identify an objective to optimize and constraints to satisfy. Finally, its output can be scrutinized to check properties and can be a posteriori completed with side constraints or additional objectives, to inject further instance-dependent information, that may have been independently learned, be available as knowledge, or as user requirements. These facilities are crucially needed in applicative contexts such as protein design (albanese2025computational, 2).
Much remains to be done around this architecture. As it has been done for SAT-Net (lim2022learning, 38), the ultimate $N(\omega)$ GM layer of our architecture could be analyzed during training to identify emerging hypothetical global properties such as symmetries or global decomposable constraints, allowing for more interpretable results.
For memory and computational efficiency reasons, we restricted ourselves to pairwise GMs, limiting the detection of complex many-body interactions. The ability of the architecture to deal with latent variables could be combined with dual/hidden representations of constraints (stergiou1999encodings, 65) to enhance its capacity to fully capture such interactions.
The language of NP-hard discrete Graphical Models with latent variables that we use as the output of our architecture is powerful. Yet, like any NP-hard modeling language, it is not necessarily optimal for concisely representing arbitrary decision NP-complete problems. To extend the range of practical problems that can be efficiently and concisely learned by our architecture, the E-PLL loss should be extended with the ability to deal with extended languages, including, e.g., clauses, linear constraints, and other global constraints and cost functions (ijcai2023p208, 10).
@@ Acknowledgements.
We thank S. Buchet and H. Bierlee for proofreading. Most of this work was done during MD’s PhD at INSA Toulouse and ANITI, INRAE. This work has been financially supported by the French national research project GMLaS under grant agreement ANR-24-CE23-3429 and has been supported by the AI-cluster ANITI. ANITI is funded by the France 2030 program under the Grant agreement ANR-23-IACL-0002.
@@ B.2. Analysing the PLL as a Fenchel-Young Loss
Fenchel-Young Losses (blondel2020learning, 12) (FYL) are losses that are built based on a regularized prediction function. In the context of SOP, when the structure $\mathcal{Y}$ is captured by a polytope and the objective function is linear, one vertex of the polytope will be an optimal solution. During learning, the learned objective parameters $\mathbf{c}$ evolve, and this leads to sudden shifts of the optimal discrete solution to a new vertex. Regularization in the convex hull $conv(\mathcal{Y})$ of the solution space $\mathcal{Y}$ provides a smoothing component that creates differentiability.
@@
For pairwise MRFs, the sufficient statistics $\phi(\mathbf{y})$ of an assignment $\mathbf{y}$ are defined over pairs of values of pairs of variables $(y_{i},y_{j})$. With the MRF negative log-likelihood (NLL) loss, also known as the Conditional Random Field (CRF) loss, the potential function is the intractable log-partition function (See (blondel2020learning, 12, p. 10)). For tractability reasons, the NPLL is sparse. Its support includes only a neighborhood $N_{1}(\mathbf{y})\subset\mathcal{Y}$ around the true output $\mathbf{y}$, obtained by changing the value of at most one variable:
@@ C.2. Robustness to Parameter $k$
Table 6. Average performances over $10$ initializations, for various values of parameter $k$. Training is up to $100$ epochs.
@@ C.3. Architecture and Training Details
For the protein design task, the neural network is composed of an input linear layer of size 128, a block to extract environment information, and a ResMLP to predict cost matrices. The environment network is composed of a gated MLP and a ResNet, repeated $6$ times (with shared weights), the output of the previous iteration being added to the input. The gatedMLP has $3$ layers, a width of $2\times 256$ and it considers the 48 nearest neighbours of each residue. All ResMLP contains $6$ layers, with a recurrent connection every 2 layers, and a width of $256$. The resulting architecture has 3.43.4M parameters.
As in most protein energy functions (alford2017rosetta, 3), amino acid pairs separated by a large distance are ignored. We use a 15 Å threshold. The neural network is trained using the Adam optimizer, with a weight decay of $10^{-3}$ and an initial learning rate of $5.10^{-4}$, divided by 10 when the validation loss decreases (with patience 0) until it reaches $10^{-*}$. A L1 regularization of $10^{-4}$ is applied to costs. To compare the NPLL loss and the E-PLL loss, we trained the same model with the same hyperparameters and starting from the same weight initialization with each of the loss. Both models run as long as the minimum LR is not reached, and therefore not necessarily for the same number of epochs.